OpenStudy (anonymous):

What is the 9th term of the geometric sequence where a1 = -5 and a6 = -5,120? a. -327,680 b. -393,216 c. -458,752 d. -524,288

OpenStudy (amistre64):

another simple formula but for geo seqs is\[a_n=a_1*r^{(n-1)}\]

OpenStudy (amistre64):

we know a6 = a1*r^5; solving for r we get: r = 5rt(a6/a1)

OpenStudy (amistre64):

using this back in to find a9 we get: a9 = a1*r^8 a9 = a1*[5rt(a6/a1)]^8

OpenStudy (amistre64):

or we can just make all that an exponent a9 = a1* (a6/a1)^(8/5)